OpenStudy (anonymous):

I will medal and fan! Find the value of csc theta for angle theta in standard position if the point at (5,-2) lies on its terminal side.

OpenStudy (anonymous):

The hypotenuse = sqrt(5^2 +(-2)^2) = sqrt(29) ....... Since the point (5,-2) lies in 4th quadrant where cosec theta is -ve Therefore cosec(theta) =-2/sqrt(29) = -2sqrt(29)/29 ....Ans(b)

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

your welcome